# Breakdown of Bulk-Radiation Correspondence in Radiative Photonic Lattices

### v2

@@ I Introduction
To better understand the bulk-radiation correspondence in radiative photonic lattices, we develop a model to describe the guided and leaky modes of photonic crystal slabs with long-range couplings and non-local responses. The model facilitates the presentation and analysis of multi-band systems. On this basis, we conduct analytical calculations under a generic multi-band configuration. The result shows that the comparability between the bulk and far-field Berry curvature is not universal; the validity is contingent upon the particular states. Further, we apply our model approach to two foundational two-dimensional lattice prototypes. Non-zero Berry curvatures and non-trivial topological charges are simultaneously found around the high symmetry points of interest. Accordingly, we separately study the correspondence on BIC and non-BIC bands. The observations obtained from numerical results are consistent with those indicated by the analytical results.
The paper is organized as follows. In Sec. II, the diffractive and radiative coupling model is introduced to construct the non-Hermitian Hamiltonian. Within this framework, we conduct step-by-step derivations to analyze a generic multi-band configuration in Sec. III. In Sec. IV and Sec. V, the model is applied to the $\Gamma$-point of a square lattice and the $K$-point of a honeycomb lattice, respectively. We comment on the main features and discuss the symmetry constraints and parameter settings in our simulations in Sec. VI before concluding in Sec. VII.
@@ II The Diffractive and Radiative Coupling Model of Photonic Lattices
The band structure for two-dimensional photonic crystal slabs can be understood by considering light cones folded at diffraction orders $(m,n)$, where $m$, $n$ are integers [46]. For a lattice with periodicity $a_{1}$ and $a_{2}$, the diffraction orders are located at multiples of the reciprocal lattice vector $\mathbf{G}=mG_{1}\mathbf{\hat{k}}_{1}+nG_{2}\mathbf{\hat{k}}_{2}$, where $G_{1,2}=2\pi/a_{1,2}$ defines the Brillouin zone. In the case of a rectangular lattice, for example, $a_{1,2}=a_{x,y}$, $\mathbf{\hat{k}}_{1,2}=\mathbf{\hat{k}}_{x,y}$ and $G_{1,2}=G_{x,y}$. We indicate the guided modes at $\mathbf{G}$ with in-plane wave vectors $\mathbf{k_{m,n}}=(k_{x}+mG_{x})\mathbf{\hat{k}}_{x}+(k_{y}+nG_{y})\mathbf{\hat{k}}_{y}$ as $|m,n\rangle$. These guided modes constitute the basis for the effective model. Here, we only consider a single "unfolded" guided mode - for instance, the fundamental TE guided mode propagating in the uncorrugated slab made of the average effective index. The thickness of the slab is chosen to ensure operation in the single-mode regime, and all other guided TE modes lie at much higher energies.
The dispersion relation of the mode $|m,n\rangle$ with a group velocity $v_{g}$ is:
$\omega_{m,n}=\omega_{0}+\frac{v_{g}}{\hbar}|\mathbf{k_{m,n}}|,$ (1)
where $\omega_{0}$ is the mode frequency at the first $\Gamma$-point. The guided modes are then coupled via a diffractive mechanism and the radiation continuum  [47, 37, 48, 49]. For TE modes, the unit polarization vector $\mathbf{\hat{p}_{m,n}}$ associated with the guided mode $|m,n\rangle$ is perpendicular to the corresponding unit in-plane wavevector $\mathbf{k_{m,n}}$. We define the angle $\theta_{m,n}$ on the $k_{x}-k_{y}$ plane as
@@
Figure 1: The geometric representation of the basis sets $\mathbf{k_{m,n}}$ and $\mathbf{E_{m,n}}$, the in-plane wavevector $\mathbf{k}$ and far-field radiation $\mathbf{E}$ are resulted from their linear combinations.
@@ III The Bulk-Radiation Correspondence under n-band Configuration
for $\alpha,\beta=L,R$. Since all four Berry curvatures carry equivalent information [51] and give rise to associated results, we restrict ourselves to $B_{RR}$ in the following discussions. The band index is omitted in the following for simplicity of notation. The complex entry for forming non-zero bulk Berry curvatures can be purely provided by the non-Hermitian part of Eq. 6, e.g., the model shown in Sec. IV. Besides, the elicitation of the bulk Berry curvature can be further promoted by having a complex coupling scheme in the Hermitian part of Eq. 6, e.g., the model shown in Sec. V.
@@ V Bulk-radiation Correspondence around the $K$-point of Honeycomb Lattices
where $\mathbf{G_{m,n}}=m\mathbf{G}_{1}+n\mathbf{G}_{2}$, and $k_{x,y}$ are relative to the $K$ point located at $K=\left(\frac{2\pi}{3a},\frac{2\pi}{3\sqrt{3}a}\right)$. Similarly as in Sec. IV, we consider coupling between these diffracted orders. A complex coupling scheme is introduced, which in essence reflects a different symmetry breaking mechanism. Further discussion is included in Sec. VI.
@@
$\hat{H}=\begin{pmatrix}\omega_{0,1}&u&u^{*}\\ u^{*}&\omega_{2,1}&u\\ u&u^{*}&\omega_{0,-1}\\ \end{pmatrix},$ (21)
where $\omega_{m,n}$ is defined in Eq. (1) with Eq. (20), while the coupling $u$ provides in-plane mode mixing between the three diffracted orders. The non-Hermitian part can be correspondingly expressed as
@@
When $u$ is a complex number, the $C_{3}$ symmetry is broken and some unique features can be captured on this platform. Particularly, net Berry curvature accumulation is found at the valleys of the non-local honeycomb lattice under the complex setting. The comparability around the $K$-point, as reported in [43], is successfully reproduced using our method, e.g., for non-BIC bands in Fig. 5. On top of that, we identify a regime where the comparability breaks down, e.g., for non-BIC bands in Fig. 6. The former case holds when the real part of $u$ is way larger than the imaginary part of $u$, while the latter case stands for the setting as the imaginary part of $u$ increases and becomes comparable to the real part of $u$.
@@ VI.1 Breakdown of Bulk-Radiation Correspondence: Physical Insights
Firstly, introducing long-range couplings, i.e., squeezing the distance between light cones, gives rise to progressive gap reduction. Thus, the models in Sec. IV and Sec. V with ultra-long-range couplings cannot be effectively treated as two-band systems. In other words, two-band regimes are not complete. Unlike the case discussed in ref. [39], the multi-band nature leads to dimensional mismatch upon projecting the bulk Hilbert space to the far-field synthetic parameter space. In general, this dimensional mismatch results in the loss of both local geometrical and global topological features of the original manifolds.
@@
Thus, the bulk-radiation correspondence is not universal in photonic lattices with unavoidable guided resonance leakage, even for those with shorter-range couplings than the cases shown in Sec. IV and Sec. V. This is consistent with the analytical results obtained in Sec. III. Additionally, we infer that the bulk-radiation correspondence also fails in non-lattice planar photonic structures with more than two decoupled degrees of freedom, where effective two-level treatment no longer applies.
@@ VI.3 Realistic Meanings of Coupling Parameter Settings
In Sec. IV, the parameters $u_{x}$ and $u_{y}$ relate to the couplings between counter-propagating guided modes along the two diagonals across the first Brillouin zone. As the couplings cross the $\Gamma$-point, the relation between $u_{x}$ and $u_{y}$ largely relies on the symmetry of the site geometry with respect to the two diagonals in the square unit cell. For instance, $u_{x}$ and $u_{y}$ are equivalent if a symmetric shape is adopted, e.g., a circle or a square oriented along $x$ and $y$, which corresponds to the cases shown in Fig. 2. In contrast, $u_{x}$ and $u_{y}$ become non-equal if an asymmetric shape is adopted, e.g. an ellipse or a rectangle, which corresponds to the cases shown in Fig. 3. The degree of asymmetry determines the $u_{x}/u_{y}$ ratio, which determines the degree of validity of the bulk radiation correspondence. Furthermore, the contrast between $u_{x,y}$ and $v$ could reflect material properties. For example, directional couplings along $k_{x}$ and $k_{y}$ become dominant in plasmonic cases. Thus, $u_{x,y}$ is supposed to be way larger than the parameter $v$. By increasing the scale difference between the parameters, we get comparable patterns as measured in ref. [41], see Fig. 8.
@@
In Sec. V, the parameter $u$ is not directly determined by the shape but by the difference between the two sites in the honeycomb unit cell. For example, $u$ is purely real if the two sites are identical. The bulk-radiation correspondence can thereby be regarded as valid for non-BIC bands regardless of the shape, which was shown in Fig. 4. However, when the two sites are not identical, e.g., different in sizes, the parameter $u$ enters the complex regime. The bulk-radiation correspondence gradually collapses as the imaginary part is increased, e.g., which corresponds to the change shown from Fig. 5 to Fig. 6. Moreover, the complex coupling brings in different AZ symmetry constraints, leading to net Berry concentration at the $K$ point (with the $K^{\prime}$ point exhibiting opposite Berry curvature accumulation). Furthermore, we reach this regime by having complex couplings in the Hermitian part of the Hamiltonian shown in Eq. 23. Thus, we predict that one can also reach this regime by employing magnetic materials and applying an external magnetic field.
@@ VII Conclusion
By developing an effective model to describe the guided and leaky modes of non-local radiative photonic lattices, we construct the bulk and the radiation Berry curvatures. We analyze the comparability between bulk and radiation topology with both analytical and numerical methods. The analytical results straightforwardly convey that the bulk-radiation correspondence is not universal, and the validity depends on the particular bulk Bloch states. The numerical results help to further exemplify specific cases where the bulk-radiation correspondence holds and collapses.
@@
The breakdown of the correspondence between far-field and bulk Berry curvature raises a question: which one dictates the system’s topological properties? While only far-field curvature is easily experimentally accessible, phenomena like wavepacket dynamics [59] and anomalous Hall velocity could offer indirect probes of the bulk Berry curvature. Comparing these indirect signatures with far-field measurements may resolve this ambiguity between bulk and radiation topology. Additionally, incorporating nonlinear effects in this model, relevant for polariton condensates and lasing, offers another direction for future studies, potentially enriching the interplay between topology, interactions, and radiative losses in non-Hermitian photonic lattices.
@@ VIII Acknowledgments
H.S. acknowledges the Icelandic Research Fund (Rannís), grant No. 239552-051. This work was partially supported by the French National Research Agency (ANR) under the project POLAROID (ANR-24-CE24-7616-01). The work in Aalto is supported by the Academy of Finland under project no. 13354165.

### v3

@@ I Introduction
To better understand the bulk-radiation correspondence in radiative photonic lattices, we develop a model to describe the guided and leaky modes of photonic crystal slabs with long-range couplings and non-local responses. The model facilitates the presentation and analysis of multi-band systems. On this basis, we conduct analytical calculations under a generic multi-band configuration. The result shows that the comparability between the bulk and far-field Berry curvature is not universal; the validity is contingent upon the particular states. Further, we apply our model to two foundational two-dimensional lattice prototypes. Non-zero Berry curvatures and non-trivial topological charges are simultaneously found around the high symmetry points of interest. Accordingly, we separately study the correspondence on BIC and non-BIC bands. The observations obtained from numerical results are consistent with those indicated by the analytical results.
The paper is organized as follows. In Sec. II, the diffractive and radiative coupling model is introduced to construct the non-Hermitian Hamiltonian. Within this framework, we conduct step-by-step derivations to analyze a generic multi-band configuration in Sec. III. In Sec. IV and Sec. V, the model is applied to the $\Gamma$-point of a square lattice and the $K$-point of a honeycomb lattice, respectively. We comment on the physical insights provided by this breakdown of bulk-radiation correspondence, followed by a detailed analysis of the symmetry constraints and parameter settings in our simulations in Sec. VI. Final conclusions are given in Sec. VII.
@@ II The Diffractive and Radiative Coupling Model of Photonic Lattices
The band structure for two-dimensional photonic crystal slabs can be understood by considering guided modes folded at diffraction orders $(m,n)$, where $m$, $n$ are integers [46]. For a lattice with periodicity $a_{1}$ and $a_{2}$, the diffraction orders are located at multiples of the reciprocal lattice vector $\mathbf{G}=mG_{1}\mathbf{\hat{k}}_{1}+nG_{2}\mathbf{\hat{k}}_{2}$, where $G_{1,2}=2\pi/a_{1,2}$ defines the Brillouin zone. In the case of a rectangular lattice, for example, $a_{1,2}=a_{x,y}$, $\mathbf{\hat{k}}_{1,2}=\mathbf{\hat{k}}_{x,y}$ and $G_{1,2}=G_{x,y}$. We indicate the guided modes in the vicinity of $\mathbf{G}$ with in-plane wave vectors $(k_{x},k_{y})\in 1^{st}BZ$ as $|m,n\rangle$. These guided modes constitute the basis for the effective model. Here, we only consider a single "unfolded" guided mode - for instance, the fundamental TE guided mode propagating in the uncorrugated slab made of the average effective index. The thickness of the slab is chosen to ensure operation in the single-mode regime, and all other guided TE modes lie at much higher energies.
The dispersion relation of the mode $|m,n\rangle$ in the vicinity of $\Gamma$-point corresponding to $\mathbf{G}$ with a group velocity $v_{g}$ is:
$\omega_{m,n}=\omega_{0}+\frac{v_{g}}{\hbar}\left(|\mathbf{k_{m,n}}|-|\mathbf{G}|\right),$ (1)
where $\omega_{0}$ is the mode frequency at the first $\Gamma$-point. The guided modes are then coupled via a diffractive mechanism and the radiation continuum  [47, 37, 48, 49]. The coefficients determining the coupling parameters can be obtained from a perturbative method based on the periodic potential profile acting on the photonic modes.
For TE modes, the unit polarization vector $\mathbf{\hat{p}_{m,n}}$ associated with the guided mode $|m,n\rangle$ is perpendicular to the corresponding unit in-plane wavevector $\mathbf{k_{m,n}}$. We define the angle $\theta_{m,n}$ on the $k_{x}-k_{y}$ plane as
@@
Figure 1: The geometric representation of the basis sets $\mathbf{k_{m,n}}$ and $\mathbf{E_{m,n}}$, the wavevector $\mathbf{k}$ of guided resonances and far-field electric field components $\mathbf{E}$ are resulted from their linear combinations.
@@ III The Bulk-Radiation Correspondence under n-band Configuration
for $\alpha,\beta=L,R$, with $\partial_{\mu,\nu}$ being different operators in k-space. Since all four Berry curvatures carry equivalent information [51] and give rise to associated results, we restrict ourselves to $B_{RR}$ in the following discussions. The band index is omitted in the following for simplicity of notation. The complex entry for forming non-zero bulk Berry curvatures can be purely provided by the non-Hermitian part of Eq. 6, e.g., the model shown in Sec. IV. Besides, the elicitation of the bulk Berry curvature can be further promoted by having a complex coupling scheme in the Hermitian part of Eq. 6, e.g., the model shown in Sec. V.
@@
The electric field components can be extracted from the texture of the Stokes parameter in k-space.
@@ V Bulk-radiation Correspondence around the $K$-point of Honeycomb Lattices
where $\mathbf{G_{m,n}}=m\mathbf{G}_{1}+n\mathbf{G}_{2}+\mathbf{K}$, and $k_{x,y}$ are relative to the $K$ point located at $\mathbf{K}=\left(\frac{2\pi}{3a},\frac{2\pi}{3\sqrt{3}a}\right)$. Similarly as in Sec. IV, we consider coupling between these diffracted orders. A complex coupling scheme is introduced, which in essence reflects a different symmetry-breaking mechanism. Further discussion is included in Sec. VI.
@@
$\hat{H}=\begin{pmatrix}\omega_{0,1}&u&t^{*}\\ u^{*}&\omega_{2,1}&u\\ t&u^{*}&\omega_{0,-1}\\ \end{pmatrix},$ (21)
where $\omega_{m,n}$ is defined in Eq. (1) with Eq. (20), while the couplings $u$, $t$ provide in-plane mode mixing between the three diffracted orders. Although $u\gtrapprox t$ for a generic setting, we keep $u=t$ since a weak perturbation on the potential profile is considered. The non-Hermitian part can be correspondingly expressed as
@@
When $u$ is a complex number, inversion symmetry is broken and some unique features can be captured on this platform. Particularly, net Berry curvature accumulation is found at the valleys of the non-local honeycomb lattice under the complex setting. The comparability around the $K$-point, as reported in [43], is successfully reproduced using our method, e.g., for non-BIC bands in Fig. 5. On top of that, we identify a regime where the comparability breaks down, e.g., for non-BIC bands in Fig. 6. The former case holds when the real part of $u$ is way larger than the imaginary part of $u$, while the latter case stands for the setting as the imaginary part of $u$ increases and becomes comparable to the real part of $u$.
@@ VI.1 Breakdown of Bulk-Radiation Correspondence: Physical Insights
Firstly, it’s not fully legitimate to ignore the multi-band nature of photonic crystal slabs with $a<\lambda$. For example, the models in Sec. IV and Sec. V cannot be effectively treated as two-band systems. The small gaps between the bands make them sense the influence from one another. In other words, two-band regimes are not complete. Unlike the case discussed in ref. [39], the multi-band nature leads to dimensional mismatch upon projecting the bulk Hilbert space to the far-field synthetic parameter space. In general, this dimensional mismatch results in the loss of both local geometrical and global topological features of the original manifolds.
@@
Thus, the bulk-radiation correspondence is not universal in photonic lattices with unavoidable guided resonance leakage. This is consistent with the analytical results obtained in Sec. III. Additionally, we infer that the bulk-radiation correspondence also fails in non-lattice planar photonic structures with more than two decoupled degrees of freedom, where effective two-level treatment no longer applies.
@@ VI.3 Realistic Meanings of Coupling Parameter Settings
In Sec. IV, the parameters $u_{x}$ and $u_{y}$ relate to the couplings between counter-propagating guided modes along the two diagonals across the first Brillouin zone. As the couplings cross the $\Gamma$-point, the relation between $u_{x}$ and $u_{y}$ largely relies on the symmetry of the site geometry with respect to the two diagonals in the square unit cell. For instance, $u_{x}$ and $u_{y}$ are equivalent if a symmetric shape is adopted, e.g., a circle or a square oriented along $x$ and $y$, which corresponds to the cases shown in Fig. 2. In contrast, $u_{x}$ and $u_{y}$ become non-equal if an asymmetric shape is adopted, e.g. an ellipse or a rectangle, which corresponds to the cases shown in Fig. 3. The degree of asymmetry determines the $u_{x}/u_{y}$ ratio, which determines the degree of validity of the bulk radiation correspondence. Furthermore, the contrast between $u_{x,y}$ and $v$ could reflect material properties. For example, directional dipolar couplings along $k_{x}$ and $k_{y}$ become dominant in plasmonic cases, where $u_{x,y}$ is supposed to be way larger than the parameter $v$. By increasing the scale difference between the parameters, we get comparable patterns as measured in ref. [41] and numerically calculated from T-matrix method in ref. [59], see Fig. 8.
@@
In Sec. V, the parameter $u$ is not directly determined by the shape but by the difference between the two sites in the honeycomb unit cell. For example, $u$ is purely real if the two sites are identical. The bulk-radiation correspondence can thereby be regarded as valid for non-BIC bands regardless of the shape, which was shown in Fig. 4. However, when the two sites are not identical, e.g., different in size, the parameter $u$ enters the complex regime. The bulk-radiation correspondence gradually collapses as the imaginary part is increased, e.g., which corresponds to the change shown from Fig. 5 to Fig. 6. Moreover, the complex coupling brings in different AZ symmetry constraints, leading to net Berry concentration at the $K$ point (with the $K^{\prime}$ point exhibiting opposite Berry curvature accumulation). Furthermore, we reach this regime by having complex couplings in the Hermitian part of the Hamiltonian shown in Eq. 23. Thus, we predict that one can also reach this regime by employing magnetic materials and applying an external magnetic field.
@@ VII Conclusion
By developing an effective model to describe the guided and leaky modes of non-local radiative photonic lattices, we construct the bulk and the radiation Berry curvatures. We analyze the comparability between bulk and radiation topology with both analytical and numerical methods. The analytical results straightforwardly convey that the bulk-radiation correspondence is not universal, and the validity depends on the particular bulk Bloch states. The numerical results further exemplify specific cases where the bulk-radiation correspondence holds and collapses.
@@
While only far-field curvature is directly measurable, the authenticity of the simulated bulk Berry curvature can be indirectly probed by phenomena like wavepacket dynamics  [60]. Comparing these indirect signatures with far-field measurements may experimentally resolve this ambiguity between bulk and radiation topology. Additionally, incorporating nonlinear effects in this model, relevant for polariton condensates and lasing, offers another direction for future studies, potentially enriching the interplay between topology, interactions, and radiative losses in non-Hermitian photonic lattices.
@@ VIII Acknowledgments
This work was partially supported by the French National Research Agency (ANR) under the project POLAROID (ANR-24-CE24-7616-01). The work in Aalto is supported by the Academy of Finland under project no. 13354165. G.S. also received support from the MUR - Italian Ministry of Research - under the Rita Levi Montalcini program. H.S. acknowledges the Icelandic Research Fund (Rannís), grant No. 239552-051 and the project No. 2022/45/P/ST3/00467 co-funded by the Polish National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339.